# RMBRec: Robust Multi-Behavior Recommendation towards Target Behaviors

### v2

@@ Abstract.
$c$s: Information systems Recommender systems
@@ 1. Introduction
Recommender systems (RS) have become essential tools of the Web ecosystem (Zhuo et al., 2024; Xu et al., 2025), helping users discover preferred items, news, or services from massive candidate pools (Chen et al., 2019; Covington et al., 2016). Traditional studies usually focus on a single type of feedback (e.g., clicks or purchases), a setting often referred to as single-behavior recommendation (Rendle et al., 2009; Chen et al., 2020; Koren et al., 2009). However, user interactions in real-world scenarios are inherently multi-dimensional, including view, collect, add-to-cart, and purchase behaviors (Jin et al., 2020; He et al., 2024b; Ma et al., 2025; Han et al., 2024). Leveraging such heterogeneous feedback, known as multi-behavior recommendation, can enrich supervision and improve the prediction of target behaviors (e.g., purchases) (Wu et al., 2024).
@@ 2. Related Works
Multi-Behavior Recommendation. Multi-behavior recommendation (MBR) enhances user modeling by incorporating heterogeneous interactions such as view, cart, and purchase (Gao et al., 2019; Jin et al., 2020). Early studies extend matrix factorization and collaborative filtering by sharing latent factors across behaviors (Gao et al., 2019). With the success of graph neural networks (GNNs) (Liu et al., 2025), models such as MBGCN (Jin et al., 2020), MB-CGCN (Cheng et al., 2023), and CRGCN (Yan et al., 2023) construct behavior-specific graphs and propagate user–item relations to capture high-order dependencies. Attention-based approaches (Yan et al., 2024; Meng et al., 2023) further introduce behavior-level attention to adaptively weigh the contribution of each behavior. While these methods effectively utilize multi-behavior signals, they usually assume that all auxiliary interactions are semantically consistent with the target behavior. However, in real-world data, auxiliary signals may be noisy or even misaligned (e.g., frequent views without purchases), which can introduce biased supervision and reduce model robustness (Gao et al., 2022; Li et al., 2023; He et al., 2024a; Liu et al., 2016).
@@
Contrastive Learning for Multi-Behavior Recommendation. To enhance robustness against noisy or insufficient supervision, recent recommendation research has introduced contrastive learning (Chuang et al., 2022; Wei et al., 2021; Cai et al., 2024b, a, 2026) that regularizes latent representations through auxiliary consistency signals. In single-behavior settings, SimGCL (Yu et al., 2022) perturbs user and item embeddings with stochastic noise and applies InfoNCE-based contrastive regularization to improve robustness and generalization. Such contrastive consistency objectives have inspired extensions to multi-behavior scenarios, where heterogeneous behaviors are explicitly modeled through graph propagation or shared latent structures. For instance, CRGCN (Yan et al., 2023) and MB-CGCN (Cheng et al., 2023) refine cross-behavior dependencies via residual or cascading propagation, while PKEF (Meng et al., 2023) adopts a parallel expert fusion structure to disentangle and balance heterogeneous behavior signals. Building on this, self-supervised approaches such as S-MBRec (Gu et al., 2022), MBSSL (Xu et al., 2023), and MISSL (Wu et al., 2024) introduce intra- and inter-behavior contrastive learning objectives to enhance semantic consistency and optimization stability. They provide auxiliary alignment signals across behaviors or interests, alleviating sparsity and improving transferability under noisy supervision. Meanwhile, BCIPM (Yan et al., 2024) incorporates behavior-contextualized preference modeling to filter irrelevant signals. Despite these advances, most of these methods regularize representations via contrastive or reconstruction objectives, yet they still assume consistent cross-behavior semantics and rarely exploit the natural heterogeneity among behaviors to achieve stable optimization and robust generalization.
@@ 4.1.1. Datasets
We evaluate our model  RMBRec  on three widely used multi-behavior datasets: Taobao [^1] https://tianchi.aliyun.com/dataset/649, TMall [^2] https://tianchi.aliyun.com/dataset/140281, and Beibei [^3] https://figshare.com/. Table 1 summarizes their statistics, including the number of users, items, interactions, and behavior type. In all datasets, purchase is regarded as the target behavior, while the others (e.g., view, add-to-cart, collect) are treated as auxiliary behaviors.
@@ 4.2. Overall Performance
We present the overall comparison results on three benchmark datasets in Table 2. Across almost all metrics, RMBRec consistently surpasses both single- and multi-behavior baselines, indicating its strong ability to utilize auxiliary signals while remaining resilient to behavioral noise and heterogeneity. The following observations can be drawn:
@@ 4.4. Robustness Study
To further examine the robustness of RMBRec under noisy environments, we conduct a controlled perturbation study on the Taobao dataset by randomly modifying auxiliary behavior interactions. Specifically, we randomly add 10%, 30%, and 50% new edges to simulate spurious or noisy behaviors, and remove 10%, 30%, and 50% existing edges to simulate incomplete or missing auxiliary information. This bidirectional perturbation allows us to evaluate the model’s resilience against both over- and under-observed behavioral signals. We compare RMBRec with three competitive multi-behavior baselines—UIPL (Yan et al., 2025), BCIPM (Yan et al., 2024), and PKEF (Meng et al., 2023)—that have demonstrated strong overall accuracy under standard settings.
As shown in Figure 3, the left plot reports the absolute performance (HR@10), while the right plot shows the relative drop ratio with respect to the noise-free condition. From the left figure, we observe that RMBRec consistently achieves the highest accuracy across all perturbation ratios, regardless of whether the noise arises from excessive or missing interactions. This demonstrates that RMBRec effectively resists the influence of random behavioral corruption and maintains reliable ranking quality even under substantial perturbations. From the right figure, although all models exhibit performance degradation as perturbations increase, the decline of RMBRec is much smaller than that of the other methods, indicating its remarkable stability. This robustness stems from the collaborative effect of the Representation Robustness Module (RRM)—which filters out semantically drifted auxiliary behaviors—and the Optimization Robustness Module (ORM)—which ensures global invariance across disturbed behavioral environments. These components prevent RMBRec from overfitting spurious patterns and enable stable recommendations under noisy or incomplete auxiliary signals in real-world settings.
@@ C.2.1. Background on IRM
Invariant Risk Minimization (IRM) (Arjovsky et al., 2019) was proposed as a principle for out-of-distribution (OOD) generalization (Liu et al., 2021; Wang et al., 2022). Its key idea is that if a representation truly captures invariant and causal factors, then a single linear classifier should perform well across multiple environments (Bai et al., 2024; Chen et al., 2025). Formally, IRM aims to find a feature extractor $\Phi$ and a classifier $w$ such that (Arjovsky et al., 2019):

### v3

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792617isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Recommender systems
@@ 1. Introduction
Recommender systems (RS) have become essential tools of the Web ecosystem (Zhuo et al., 2024; Xu et al., 2025; Cheng et al., 2018), helping users discover preferred items, news, or services from massive candidate pools (Chen et al., 2019; Covington et al., 2016). Traditional studies usually focus on a single type of feedback (e.g., clicks or purchases), a setting often referred to as single-behavior recommendation (Rendle et al., 2009; Chen et al., 2020; Koren et al., 2009). However, user interactions in real-world scenarios are inherently multi-dimensional, including view, collect, add-to-cart, and purchase behaviors (Jin et al., 2020; He et al., 2024b; Ma et al., 2025; Han et al., 2024). Leveraging such heterogeneous feedback, known as multi-behavior recommendation, can enrich supervision and improve the prediction of target behaviors (e.g., purchases) (Wu et al., 2024).
@@ 2. Related Works
Multi-Behavior Recommendation. Multi-behavior recommendation (MBR) models user preferences by leveraging heterogeneous interactions such as view, cart, and purchase (Gao et al., 2019; Jin et al., 2020). Early methods extend matrix factorization or collaborative filtering by sharing latent factors across behaviors (Gao et al., 2019). With the advancement of graph neural networks (GNNs), subsequent studies construct behavior-specific graphs and perform cross-behavior propagation to capture high-order user–item dependencies (Jin et al., 2020; Cheng et al., 2023; Yan et al., 2023; Liu et al., 2021a). Attention-based approaches further adaptively weigh different behaviors to enhance flexibility (Yan et al., 2024; Meng et al., 2023). Despite their effectiveness, most existing MBR methods implicitly assume that auxiliary behaviors are semantically consistent with the target behavior. In practice, however, auxiliary interactions are often noisy or weakly correlated with target preferences (e.g., frequent views without purchases), which may introduce biased supervision and degrade robustness (Gao et al., 2022; Li et al., 2023; He et al., 2024a; Liu et al., 2016; Li et al., 2025).
@@
Contrastive Learning for Multi-Behavior Recommendation. To enhance robustness against noisy or insufficient supervision, recent recommendation research has widely adopted contrastive learning to regularize latent representations through auxiliary consistency signals (Chuang et al., 2022; Wei et al., 2021; Cai et al., 2024b, a, 2026). In single-behavior settings, SimGCL (Yu et al., 2022) improves robustness and generalization by perturbing user and item embeddings with stochastic noise and enforcing InfoNCE-based contrastive consistency. Building on this paradigm, contrastive learning has been extended to multi-behavior recommendation, where heterogeneous behaviors are modeled via graph propagation or shared latent structures. Representative methods such as CRGCN (Yan et al., 2023) and MB-CGCN (Cheng et al., 2023) refine cross-behavior dependencies through cascading or residual propagation, while PKEF (Meng et al., 2023) adopts a parallel expert fusion framework to disentangle and balance heterogeneous behavior signals. Furthermore, self-supervised approaches including S-MBRec (Gu et al., 2022), MBSSL (Xu et al., 2023), and MISSL (Wu et al., 2024) introduce intra- and inter-behavior contrastive objectives to enhance semantic consistency and optimization stability under noisy supervision. Meanwhile, BCIPM (Yan et al., 2024) incorporates behavior-contextualized preference modeling to filter task-irrelevant signals. Despite these advances, most existing methods still regularize representations via contrastive or reconstruction objectives, implicitly assuming consistent cross-behavior semantics, and thus remain limited in exploiting intrinsic behavioral heterogeneity to address behavior-induced optimization instability and achieve robust generalization.
@@ 4.1.1. Datasets
We evaluate our model  RMBRec  on three widely used multi-behavior datasets: Taobao [^1] https://tianchi.aliyun.com/dataset/649, TMall [^2] https://tianchi.aliyun.com/dataset/140281, and Beibei. Table 1 summarizes their statistics, including the number of users, items, interactions, and behavior type. In all datasets, purchase is regarded as the target behavior, while the others (e.g., view, add-to-cart, collect) are treated as auxiliary behaviors.
@@ 4.2. Overall Performance
We present the overall comparison results on three benchmark datasets in Table 2. Across almost all metrics, RMBRec consistently surpasses both single- and multi-behavior baselines, indicating its strong ability to utilize auxiliary signals while remaining resilient to behavioral noise and heterogeneity.
@@ 4.4. Robustness Study
To further examine the robustness of RMBRec under noisy environments, we conduct a controlled perturbation study on the Taobao dataset by randomly modifying auxiliary behavior interactions. Specifically, we randomly add 10%, 30%, and 50% new edges to simulate spurious behaviors and remove 10%, 30%, and 50% existing edges to simulate missing auxiliary information, enabling evaluation under both over- and under-observed conditions. We compare RMBRec with three competitive multi-behavior baselines—UIPL (Yan et al., 2025), BCIPM (Yan et al., 2024), and PKEF (Meng et al., 2023)—that perform strongly under standard settings.
As shown in Figure 3, the left plot reports absolute performance (HR@10), while the right plot shows the relative drop ratio compared to the noise-free condition.  RMBRec consistently achieves the best performance across all perturbation ratios, regardless of whether noise is introduced by excessive or missing interactions, and exhibits a substantially smaller performance decline than competing methods. This robustness arises from the collaborative effect of the Representation Robustness Module (RRM), which suppresses semantically drifted auxiliary behaviors, and the Optimization Robustness Module (ORM), which enforces global invariance across disturbed behavioral environments, enabling stable and reliable recommendation performance under severe noise.
@@ C.2.1. Background on IRM
Invariant Risk Minimization (IRM) (Arjovsky et al., 2019) was proposed as a principle for out-of-distribution (OOD) generalization (Liu et al., 2021b; Wang et al., 2022). Its key idea is that if a representation truly captures invariant and causal factors, then a single linear classifier should perform well across multiple environments (Bai et al., 2024; Chen et al., 2025). Formally, IRM aims to find a feature extractor $\Phi$ and a classifier $w$ such that (Arjovsky et al., 2019):